OpenStudy (anonymous):

find the angles between the lines: y=2x-6 & 2x+3y=18

OpenStudy (amistre64):

we have 2 vectors that make up the slope

OpenStudy (amistre64):

a=<1, 2>; |a| = sqrt(5) b=<3,-2>; |b| = sqrt(13) ---------- 3-4 = -1 = a.b cos(t) = a.b / |a||b| t = cos-1 (-1/sqrt{5*13}) t = cos-1 (-1/sqrt{65}) and the other one should be pi - t

OpenStudy (amistre64):

if im lucky :) we get about 97 and 83

OpenStudy (anonymous):

yeah the answer is 97.1 and 82.9. Can you explain how you did that?

OpenStudy (amistre64):

.....i did lol

OpenStudy (amistre64):

but i used vector concepts that you prolly aint used to

OpenStudy (anonymous):

right

OpenStudy (amistre64):

we can draw it out and get the same idea i spose

OpenStudy (anonymous):

how do you know the coordinates i.e. 1,2 3,-2

OpenStudy (amistre64):



OpenStudy (amistre64):

slopes are vectors; y/x is a slope so: <x,y> is the vector y/x = 2/1 = <1,2> y/x = -2/3 = <3,-2>

OpenStudy (amistre64):

but despite that; we we draw the lines and slope the we can get the tangents needed here tan(a) = 2/1 a = tan-1 (2/1) = 1.1071 tan(b) = 2/3 b = tan-1 (2/3) = .588 c = pi - 1.1071 - .588 = 1.4465 radians * 180/pi = 82.8 180 - 82.8 ------- 97.2

OpenStudy (anonymous):

i don't understand greek but thanks

OpenStudy (amistre64):

:)

OpenStudy (amistre64):

if your material that your working with has a technique or method; we could try to work that out

OpenStudy (amistre64):

but just trying to solve this on the fly i have to resort to my own ingenuity

OpenStudy (anonymous):

which i understand... but you're using variable which have no significance to me like pi and c... i haven't learned it in school im just trying to get ahead... I have another question though...

OpenStudy (anonymous):

for finding the equation of a line such as (-2,4) and (1,2) could you explain how to find b(y-int)?

OpenStudy (amistre64):

the most important think about a line are, well, 2 things; slope, and a point to anchor it to

OpenStudy (amistre64):

you have 2 points given so we need to determine a slope

OpenStudy (anonymous):

okay so delta y/delta x right?

OpenStudy (amistre64):

\[slope=\frac{1-Py}{x-Px}\] which means we subtract one point from the other and stack y/x

OpenStudy (amistre64):

y - Py ....

OpenStudy (amistre64):

this would be ideal if we moved both point so that one of them was at the origin (0,0) (-2,4) (1,2) +2-4 +2-4 ------ ------ (0,0) (3,-2) now we are able to see the slope as -2/3 do you know why?

OpenStudy (anonymous):

no that was my next question lol

OpenStudy (amistre64):

from 0,0 to 3,-2 ; what is our change in y? what is our change in x?

OpenStudy (anonymous):

delta x = 3 and delta y= -2

OpenStudy (amistre64):

correct; so deltay/deltax as you see it is then: -2/3

OpenStudy (amistre64):

do you understand how I was able to get 0,0 and 3,-2 to work from?

OpenStudy (anonymous):

yeah because either of the points will yield the same answer

OpenStudy (amistre64):

correct, and we are able to move them to a more convient position to measure things from graphically; formulawise its just subtracting points

OpenStudy (amistre64):

with this slope; we need to reform our slope equation but with 1 point and a generic point

OpenStudy (anonymous):

okay so we can say y=(-2/3)+b as in y=mx+b therefore b= 8/3?

OpenStudy (anonymous):

*y=(-2/3)x+b

OpenStudy (amistre64):

perhaps, lets see where it leads :) ill use the point(1,2) and have a generic (x,y) \[slope=\frac{y-Py}{x-Px}\] \[\frac{-2}{3}=\frac{y-1}{x-2}\] \[(x-2)*\frac{-2}{3}=\frac{y-1}{\cancel{x-2}}*\cancel{(x-2)}\] \[\frac{-2}{3}(x-2)=y-1\] \[\frac{-2}{3}x-\frac{2}{3}(-2)=y-1\] \[\frac{-2}{3}x+\frac{4}{3}=y-1\] \[\frac{-2}{3}x+\frac{4}{3}+1=y\cancel{-1+1}^{ 0}\] \[\frac{-2}{3}x+\frac{4}{3}+1=y\] \[\frac{-2}{3}x+\frac{4}{3}+1*\frac{3}{3}=y\] \[\frac{-2}{3}x+\frac{4}{3}+\frac{3}{3}=y\] \[\frac{-2}{3}x+\frac{4+3}{3}=y\] \[\frac{-2}{3}x+\frac{7}{3}=y\] I would say from this that b = 7/3

OpenStudy (amistre64):

there are shortcuts you can take; but this IS the details that you should be familiar with

OpenStudy (anonymous):

okay i'll look this over thanks for your help

OpenStudy (amistre64):

yw

OpenStudy (anonymous):

http://www.purplemath.com/modules/strtlneq.htm this says that b=8/3 rather than 7/3?